# Mixing time of a matrix random walk generated by elementary transvections

### v1

@@ Abstract.
We consider a Markov chain on invertible $n\times n$ matrices with entries in $\mathbb{Z}_{2}$ which moves by picking an ordered pair of distinct rows and add the first one to the other, modulo $2$. We establish a logarithmic Sobolev inequality with constant $n^{2}$, which yields an upper bound of $O(n^{2}\log n)$ on the mixing time. We also study another chain which picks a row at random and randomize it conditionally on the resulting matrix being invertible, and show that the mixing time is between $n\log n$ and $4n\log n$.
@@ 1. Setting and main result
We consider a random walk on $\Omega=\mathrm{SL}_{n}(\mathbb{Z}_{2})$, the set of invertible binary matrices, generated by elementary transvections: at each time, the walk proceeds by picking two distinct rows at random, and adding one to the other, modulo $2$. Letting $E_{i,j}$ be the $n\times n$ matrix with a $1$ in position $(i,j)$, and 00 elsewhere, the transition matrix $P$ of the walk is defined by
@@
Since the elementary transvections $I+E_{i,j}$ generate $\Omega$, the chain is irreducible. By symmetry of the matrix $P$, the uniform distribution $\pi$ over $\Omega$ is the unique stationary distribution. Since the chain is also aperiodic (see [6, Proposition 4.4]), the chain converges to $\pi$ as time tends to infinity. To quantify the speed of this convergence, one usually considers the total variation mixing time defined, for all $\varepsilon\in(0,1)$, by
$t_{\mathsf{mix}}(P,\varepsilon)\;=\;\min\{t\geq 0\,,\,\max_{x\in\Omega}\|P^{t}(x,\cdot)-\pi\|_{\mathsf{tv}}\leq\varepsilon\}\,,$
where $\|\mu-\nu\|_{\mathsf{tv}}=\max_{A\subset\Omega}|\mu(A)-\nu(A)|$ is the total-variation distance.
@@
There exists an absolute constant $C>0$ such that for all $\varepsilon>0$, the mixing time of the chain $P$ satisfies
$t_{\mathsf{mix}}(P,\varepsilon)\;\leq\;Cn^{2}\log\left(\frac{n}{\varepsilon}\right)\,.$
Theorem 1 comes from an upper bound of $O(n^{2})$ on the logarithmic Sobolev constant $\mathrm{C}_{\mathsf{LS}}(P)$ of the chain, defined as the smallest constant $c>0$ such that for all functions $f\colon\Omega\to\mathbb{R}$, we have
@@
${\mathcal{E}}_{P}(f,f)\;=\;\sum_{x,y\in\Omega}\pi(x)P(x,y)\left(f(x)-f(y)\right)f(x)\,.$
A classical result is that
$t_{\mathsf{mix}}(P,\varepsilon)\;\leq\;\left\lceil\mathrm{C}_{\mathsf{LS}}(P)\left(\log\log\left(\frac{1}{\pi_{\star}}\right)+2\log\left(\frac{1}{\varepsilon}\right)\right)\right\rceil\,,$
where $\pi_{\star}=\min_{x\in\Omega}\pi(x)$ (see for instance [5] and [11]). Since in our case we have $\pi_{\star}\asymp\frac{1}{2^{n^{2}}}$, we see that Theorem 1 will follow from the following proposition.
@@
$\mathrm{C}_{\mathsf{LS}}(P)\;\lesssim\;n^{2}\,.$
@@
Interestingly, a crucial ingredient in the proof of Proposition 2 is the bound of $O(n)$ on the Poincaré constant (inverse of the spectral gap), obtained by Kassabov [9].
Remark 1.
Proposition 2 even yields an upper bound of $O(n^{2}\log n)$ for the $\ell^{2}$-mixing time, at least for the continuous-time and lazy versions of the chain.
We also study another chain on $\Omega$, whose kernel is denoted by $Q$, and whose dynamics is as follows: if the current state is a matrix $x\in\Omega$ with rows $\ell_{1},\dots,\ell_{n}$, then an index $i$ between $1$ and $n$ is chosen uniformly at random, and row $\ell_{i}$ is replaced by $\ell_{i}+\sum_{j\neq i}a_{j}\ell_{j}$, wit $(a_{j})_{j\neq i}$ independent Bernoulli random variables with parameter $1/2$. In other words, the chain $Q$ randomizes a uniformly chosen row, conditionally on the event that the resulting matrix is invertible (it is the Gibbs sampler for the joint distributions of rows). The following proposition shows that the mixing time of $Q$ lies between $n\log n$ and $4n\log n$ whatever the value of the threshold $\varepsilon$, establishing a pre-cutoff.
Proposition 3.
For all $\varepsilon\in(0,1)$, there exists a constant $c_{\varepsilon}>0$ such that
[…]
In particular, the chain $Q$ has a pre-cutoff.
@@ 2. Related work
To our knowledge, the chain $P$ was first studied by Diaconis and Saloff-Coste [6], who used a comparison with another chain on $\Omega$ studied by Hildebrand [8] to get an upper bound of $O(n^{2})$ on the relaxation time (the inverse of the spectral gap), implying an upper bound of $O(n^{4})$ on the $\ell^{2}$-mixing time. Then the results of Kassabov [9] on the Kazhdan constant yield an upper bound of $O(n^{3})$ on the $\ell^{2}$-mixing time, which was the best known upper bound for total variation mixing as well. As for the lower bound, a simple counting argument shows that the total-variation mixing time can be lower bounded by $\Omega\left(\tfrac{n^{2}}{\log n}\right)$ (which is actually an estimate of the diameter of the underlying graph, see Andrén et al. [1], Christofides [4]). It would be interesting to know whether a lower bound of $\Omega(n^{2}\log n)$ holds, and more ambitiously, whether this chain exhibits the cutoff phenomenon.
@@
The protocol proposed by Sotiraki [15] works as follows: starting with the identity matrix in $\text{SL}_{n}(\mathbb{Z}_{2})$, the prover runs the chain with kernel $P$ for a specific duration $t$. The resulting matrix $A_{t}$ is made public (acting as the public key), but only the prover knows the chain’s trajectory. Then, when he needs to authenticate, the prover requests a binary vector of length $n$ from the verifier, and the challenge is to compute $y=A_{t}x$ quickly. Since the prover is aware of the chain’s trajectory, he can apply the same row operations to $x$ that were used to generate $A_{t}$, thus providing the correct answer in time $t$. On the other hand, if $t$ is sufficiently large, a dishonest party may struggle to distinguish $A_{t}$ from a matrix chosen uniformly at random within polynomial time (indicating that the chain is computationally mixed). In this case, the dishonest party’s best option would be to perform regular matrix-vector multiplication, which typically requires $n^{2}$ operations. Therefore, if the prover selects $t$ as the computational mixing time, and if $t$ is proven to be significantly smaller than $n^{2}$, the verifier can successfully distinguish between honest and dishonest parties. We are not aware on quantitative bounds designed specifically for the computational mixing time (when only test functions that are computable in polynomial time are allowed) but the present paper shows that it is $O(n^{2}\log n)$, which supports the validity of the descirbed authentication protocol.
One may notice that the restricted dynamics on the first $k\leq n$ columns of the matrix is still a Markov chain. For $k=1$, Ben-Hamou and Peres [2] established a cutoff at time $\tfrac{3}{2}n\log n$ with a window of order $n$. Later, the results of Peres, Tanaka and Zhai [14] on finite groups allowed to generalize this cutoff for a fixed number $k$ of columns.
Another related chain is the random walk on upper triangular matrices with entries in $\mathbb{Z}_{q}$ (where $q\geq 2$ is an integer) and ones along the diagonal, where at each step a uniformly chosen row is added to the row above. For $q=2$, Peres and Sly [13] showed that the mixing time is $O(n^{2})$, and Nestoridi and Sly [12] then obtained optimal bounds for the mixing time both in $n$ and $q$ (we refer the reader to this last paper for more reference on this process).
@@ 3. Proof of Proposition 2
$\forall x\in\Omega^{*}\,,\;g(x)\;=\;\begin{cases}f(x)&\text{ if $x\in\Omega$,}\\ \mathbb{E}_{\pi}[f]&\text{ if $x\not\in\Omega$.}\end{cases}$
Denoting by $\mu$ the uniform measure on $\Omega^{*}$, we have $\mu(x)=2^{-n^{2}}$ for all $x\in\Omega^{*}$, and
@@
Using the variational characterization of entropy and the fact that $t^{2}\log\left(\frac{t^{2}}{s^{2}}\right)-t^{2}+s^{2}\geq 0$ for all $s,t\in\mathbb{R}$, we have
$\operatorname{Ent}_{\pi}(f^{2})$ $=$ $\inf_{u\in\mathbb{R}}\sum_{x\in\Omega}\pi(x)\left(f^{2}(x)\log\left(\frac{f^{2}(x)}{u}\right)-f^{2}(x)+u\right)$
$\leq$ $\frac{1}{\mu(\Omega)}\inf_{u\in\mathbb{R}}\sum_{x\in\Omega^{*}}\mu(x)\left(g^{2}(x)\log\left(\frac{g^{2}(x)}{u}\right)-g^{2}(x)+u\right)$
$\leq$ $\frac{1}{\mu(\Omega)}\sum_{x\in\Omega^{*}}\mu(x)\left(g^{2}(x)\log\left(\frac{g^{2}(x)}{\mathbb{E}_{\mu}[g^{2}]}\right)-g^{2}(x)+\mathbb{E}_{\mu}[g^{2}]\right)$
$=$ $\frac{1}{\mu(\Omega)}\operatorname{Ent}_{\mu}(g^{2})\,.$
Under the measure $\mu$, the lines $L_{1},\dots,L_{n}$ of the matrix are independent, so by the sub-additivity of entropy for product measures, we have
$\operatorname{Ent}_{\mu}(g^{2})$$\leq$ $\sum_{i=1}^{n}\mathbb{E}_{\mu}\left[\operatorname{Ent}_{\mu}\left(g^{2}\,|\,(L_{j})_{j\neq i}\right)\right]$
For each $i\in[n]$, conditionally on $(L_{j})_{j\neq i}=(\ell_{j})_{j\neq i}$, observe that if the family $(\ell_{j})_{j\neq i}$ does not span a subset of dimension $n-1$, then, whatever the value of $L_{i}$, the function $g$ is constant equal to $\mathbb{E}_{\pi}[f]$, so the conditional entropy is zero. Otherwise, if the family $(\ell_{j})_{j\neq i}$ spans a subset of dimension $n-1$, then one may complete it with some vector $\ell_{i}$ to form a basis of $\{0,1\}^{n}$. The conditional law of $L_{i}$ is then that of $\sum_{j=1}^{n}a_{j}\ell_{j}$, where $(a_{j})_{j=1}^{n}$ is uniformly distributed over $\{0,1\}^{n}$. Using the logarithmic Sobolev inequality for the hypercube, we have
$\operatorname{Ent}_{\mu}\left(g^{2}\,|\,(L_{j})_{j\neq i}=(\ell_{j})_{j\neq i}\right)\leq\frac{1}{2}\sum_{j=1}^{n}\mathbb{E}_{\mu}\left[\left(g(X)-g(X^{i\leftarrow j})\right)^{2}\,\big{|}\,(L_{j})_{j\neq i}=(\ell_{j})_{j\neq i}\right]\,,$
@@
One may now conclude with the powerful results of Kassabov [9, Section 3.2] which yield
$\operatorname{Var}_{\pi}(f)\leq 4(31\sqrt{n}+700)^{2}{\mathcal{E}}_{P}(f,f)\,.$
4. Proof of Proposition 3
In this section, we study the chain $Q$ which at each step, picks an index $i$ uniformly at random, and replace row $\ell_{i}$ by $\ell_{i}+\sum_{j\neq i}a_{j}\ell_{j}$, with $(a_{j})_{j\neq i}$ independent Bernoulli random variables with parameter $1/2$.
Let us start with the upper bound on the mixing time, which will be obtained with the path coupling method.
Let $G=(\Omega,E)$ be the graph induced naturally by the kernel $P$, i.e.
$E\;=\;\left\{\{x,y\}\subset\Omega\,,\;P(x,y)>0\right\}\,,$
and let $\rho$ be the graph distance on $G$. Let $x$ and $y$ be two neighbors in $G$. This means that, if $\ell_{1},\dots,\ell_{n}$ are the rows of $x$ and $r_{1},\dots,r_{n}$ the rows of $y$, then there exist $1\leq i\neq j\leq n$ such that $r_{j}=\ell_{j}+\ell_{i}$ and, for all $k\neq j$, $r_{k}=\ell_{k}$. We couple the distributions $Q(x,\cdot)$ and $Q(y,\cdot)$ by choosing the same uniform index $I$ in both chains for the row to be randomized and, if $(a_{j})_{j\neq I}$ (respectively $(b_{j})_{j\neq I}$) are the Bernoulli random variables used to update the chain started at $x$ (respectively at $y$), then those two vectors are coupled as follows:
  1. (1)
    if $I=j$, then we set $b_{i}=1-a_{i}$ and $b_{k}=a_{k}$ for all $k\neq i,j$;
  2. (2)
    if $I=i$, then we set $b_{k}=a_{k}$ for all $k\neq i$;
  3. (3)
    if $I=\ell\neq i,j$, then we set $b_{i}=a_{i}+a_{j}$ (modulo $2$) and $b_{k}=a_{k}$ for all $k\neq\ell,i$.
Observe that in situation $(1)$, which occurs with probability $1/n$, the distance $\rho$ decreases from $1$ to 00, while in situation $(2)$, which also occurs with probability $1/n$, the distance either increases by $1$ with probability $1/2$ if $a_{j}=1$, or stays the same if $a_{j}=0$. In situation $(3)$, the distance always stays the same. Hence, if $(X_{1},Y_{1})$ is generated according to this coupling, we have
[…]
By the path coupling method of [3] (see also [10, Theorem 14.6]), this implies that for any probability distributions $\mu$ and $\nu$ on $\Omega$,
$W(\mu Q,\nu Q)\;\leq\;\left(1-\frac{1}{2n}\right)W(\mu,\nu)\,,$
where $W(\mu,\nu)=\min_{X\sim\mu,Y\sim\nu}\mathbb{E}[\rho(X,Y)]$ is the Wasserstein distance induced by the metric $\rho$. This in turn implies that for all $\varepsilon\in(0,1)$, we have
$t_{\mathsf{mix}}(Q,\varepsilon)\;\leq\;\left\lceil 2n\left(\log\left(\textsf{diam}(G)\right)+\log\left(\frac{1}{\varepsilon}\right)\right)\right\rceil\,,$
where $\textsf{diam}(\Omega)$ is the diameter of the graph $G$, see [10, Corollary 14.8]. By Andrén et al. [1], we have
$\textsf{diam}(G)\;=\;O\left(\frac{n^{2}}{\log n}\right)\,.$
Hence, there exists $c_{\varepsilon}>0$ such that
[…]
which concludes the proof of the lower bound.
For the lower bound, we use a standard coupon collector approach: if the chain is started at the identity matrix $I$ and if there exists some row that has not been chosen for randomization, then the matrix has a row of $n-1$ zeros, an event which has probability $O(n2^{-n})$ under the stationary distribution. More precisely, letting $A$ be the event that there exists a row with $n-1$ zeros and $\tau$ the first time when all rows have been touched, we have
$Q^{t}(I,A)\geq\mathbb{P}(\tau>t)\,.$
Using the fact that the indicator variables
$I_{t}(i)={\mathbf{1}}_{\left\{\text{row $i$ has been touched before time $t$}\right\}}\,,\;i=1,\dots n$
are negatively associated (see for instance [7]), we have
$\mathbb{P}(\tau\leq t)=\mathbb{P}\left(\bigcap_{i=1}^{n}\{I_{t}(i)=1\}\right)\leq\prod_{i=1}^{n}\mathbb{P}(I_{t}(i)=1)=\left(1-\left(1-\frac{1}{n}\right)^{t}\right)^{n}\,,$
which is less that $1-\varepsilon$ for $t=n\log n-\log\log((1-\varepsilon)^{-1})n+o(n)$. For such time $t$, we thus have
$\|Q^{t}(I,\cdot)-\pi\|_{\mathsf{tv}}\geq Q^{t}(I,A)-\pi(A)\geq\varepsilon+o(1)\,,$
which concludes the proof of the lower bound.

### v2

@@ Abstract.
We consider a Markov chain on invertible $n\times n$ matrices with entries in $\mathbb{Z}_{2}$ which moves by picking an ordered pair of distinct rows and add the first one to the other, modulo $2$. We establish a logarithmic Sobolev inequality with constant $n^{2}$, which yields an upper bound of $O(n^{2}\log n)$ on the mixing time.
@@ 1. Setting and main result
We consider a random walk on $\Omega=\mathrm{SL}_{n}(\mathbb{Z}_{2})$, the set of invertible binary matrices, generated by elementary transvections: at each time, the walk proceeds by picking two distinct rows at random, and adding one to the other, modulo $2$. Letting $E_{i,j}$ be the $n\times n$ matrix with a $1$ in position $(i,j)$, and 0 elsewhere, the transition matrix $P$ of the walk is defined by
@@
Since the elementary transvections $I+E_{i,j}$ generate $\Omega$, the chain is irreducible. By symmetry of the matrix $P$, the uniform distribution $\pi$ over $\Omega$ is the unique stationary distribution. Let us point out here that $\Omega$ is a significant subset of the set of all binary matrices. Indeed
$\frac{|\Omega|}{2^{n^{2}}}\;=\;\prod_{k=1}^{n}\left(1-2^{-k}\right)\;\approx\;0.288788\,.$
Since the chain is also aperiodic (see [5, Proposition 4.4]), the chain converges to $\pi$ as time tends to infinity. To quantify the speed of this convergence, one usually considers the total variation mixing time defined, for all $\varepsilon\in(0,1)$, by
$t_{\mathsf{mix}}(\varepsilon)\;=\;\min\{t\geq 0\,,\,\max_{x\in\Omega}\|P^{t}(x,\cdot)-\pi\|_{\mathsf{tv}}\leq\varepsilon\}\,$,
where $\|\mu-\nu\|_{\mathsf{tv}}=\max_{A\subset\Omega}|\mu(A)-\nu(A)|$ is the total-variation distance. A stronger notion is the $\ell_{2}$-mixing time defined, for all $\varepsilon\in(0,1)$, by
$t_{\mathsf{mix}}^{(2)}(\varepsilon)\;=\;\min\{t\geq 0\,,\,\max_{x\in\Omega}\left\|\frac{P^{t}(x,\cdot)}{\pi(\cdot)}-1\right\|_{2,\pi}\leq\varepsilon\}\,$,
where $\|f\|_{2,\pi}=\left(\sum_{\in\Omega}\pi(x)f(x)^{2}\right)^{1/2}$ is the $\ell_{2}$-norm with respect to $\pi$. By Cauchy-Schwarz inequality, we have
$t_{\mathsf{mix}}(\varepsilon)\;\leq\;t_{\mathsf{mix}}^{(2)}(2\varepsilon)\,.$
@@
There exists absolute constants $A>0$ and $B>0$ such that for all $\varepsilon>0$, the $\ell_{2}$-mixing time of the chain $P$ satisfies
$t_{\mathsf{mix}}^{(2)}(\varepsilon)\;\leq\;An^{2}\log n+Bn\log\left(\frac{1}{\varepsilon}\right)\,.$
Theorem 1 comes from an upper bound of $O(n^{2})$ on the logarithmic Sobolev constant $\mathrm{C}_{\mathsf{LS}}$ of the chain, defined as the smallest constant $c>0$ such that for all functions $f\colon\Omega\to\mathbb{R}$, we have
@@
${\mathcal{E}}_{P}(f,f)\;=\;\frac{1}{2}\sum_{x,y\in\Omega}\pi(x)P(x,y)\left(f(x)-f(y)\right)^{2}\,.$
By [4, Corollary 3.8], we have
$t_{\mathsf{mix}}^{(2)}(\varepsilon)\;\leq\;\frac{\mathrm{C}_{\mathsf{LS}}}{4}\log\log\left(\frac{1}{\pi_{\star}}\right)+\lambda_{*}^{-1}\log\left(\frac{\sqrt{1+2e^{2}}}{\varepsilon}\right)+1\,$,
where $\pi_{\star}=\min_{x\in\Omega}\pi(x)$ and $\lambda_{*}$ is the absolute spectral gap of $P$. By [5, Proposition 4.4] and the powerful results of [7, Section 3.2], we have $\lambda_{*}^{-1}=O(n)$. Finally, since $\pi_{\star}\asymp\frac{1}{2^{n^{2}}}$, we see that Theorem 1 will follow from the following proposition.
@@
$\mathrm{C}_{\mathsf{LS}}\;\lesssim\;n^{2}\,.$
@@
Interestingly, a crucial ingredient in the proof of Proposition 2 is the bound of $O(n)$ on the Poincaré constant (inverse of the spectral gap), obtained by [7].
@@ 2. Related work
To our knowledge, the chain $P$ was first studied by [5], who used a comparison with another chain on $\Omega$ studied by [6] to get an upper bound of $O(n^{2})$ on the relaxation time (the inverse of the spectral gap), implying an upper bound of $O(n^{4})$ on the $\ell^{2}$-mixing time. Then the results of [7] on the Kazhdan constant yield an upper bound of $O(n)$ on the relaxation time (inverse spectral gap), and thus of $O(n^{3})$ on the $\ell^{2}$-mixing time, which was the best known upper bound for total variation mixing as well. As for the lower bound, a simple counting argument shows that the total-variation mixing time can be lower bounded by $\Omega\left(\tfrac{n^{2}}{\log n}\right)$ (which is actually an estimate of the diameter of the underlying graph, see [1, 3]). It would be interesting to know whether a lower bound of $\Omega(n^{2}\log n)$ holds, and more ambitiously, whether this chain exhibits the cutoff phenomenon.
@@
The protocol proposed by [11] works as follows: starting with the identity matrix in $\text{SL}_{n}(\mathbb{Z}_{2})$, the prover runs the chain with kernel $P$ for a specific duration $t$. The resulting matrix $A_{t}$ is made public (acting as the public key), but only the prover knows the chain’s trajectory. Then, when he needs to authenticate, the prover requests a binary vector of length $n$ from the verifier, and the challenge is to compute $y=A_{t}x$ quickly. Since the prover is aware of the chain’s trajectory, he can apply the same row operations to $x$ that were used to generate $A_{t}$, thus providing the correct answer in time $t$. On the other hand, if $t$ is sufficiently large, a dishonest party may struggle to distinguish $A_{t}$ from a matrix chosen uniformly at random within polynomial time (indicating that the chain is computationally mixed). In this case, the dishonest party’s best option would be to perform regular matrix-vector multiplication, which typically requires $n^{2}$ operations. Therefore, if the prover selects $t$ as the computational mixing time, and if $t$ is proven to be significantly smaller than $n^{2}$, the verifier can successfully distinguish between honest and dishonest parties. Given the present result on the total variation mixing time $O(n^{2}\log n)$, finding the computational mixing time (when only test functions that are computable in polynomial time are allowed) would be very interesting in this context.
One may notice that the restricted dynamics on the first $k\leq n$ columns of the matrix is still a Markov chain. For $k=1$, [2] established a cutoff at time $\tfrac{3}{2}n\log n$ with a window of order $n$. Later, the results of [10] on finite groups allowed to generalize this cutoff for a fixed number $k$ of columns.
Another related chain is the random walk on upper triangular matrices with entries in $\mathbb{Z}_{q}$ (where $q\geq 2$ is an integer) and ones along the diagonal, where at each step a uniformly chosen row is added to the row above. For $q=2$, [9] showed that the mixing time is $O(n^{2})$, and [8] then obtained optimal bounds for the mixing time both in $n$ and $q$ (we refer the reader to this last paper for more reference on this process).
@@ 3. Proof of Proposition 2
By the variational characterization of entropy, we have
$\operatorname{Ent}_{\pi}(f^{2})$ $=$ $\inf_{u\in\mathbb{R}_{+}}\sum_{x\in\Omega}\frac{1}{|\Omega|}\left(f^{2}(x)\log\left(\frac{f^{2}(x)}{u}\right)-f^{2}(x)+u\right)\,.$
Using the fact that $t\log\left(\frac{t}{u}\right)-t+u\geq 0$ for all $t,u\in\mathbb{R}_{+}$, one may extend the sum to $\Omega^{*}$ and replace $f$ by $g$ (which coincide with $f$ on $\Omega$) without increasing the value. More precisely, we have
$\operatorname{Ent}_{\pi}(f^{2})$ $\leq$ $\inf_{u\in\mathbb{R}_{+}}\sum_{x\in\Omega^{*}}\frac{1}{|\Omega|}\left(g^{2}(x)\log\left(\frac{g^{2}(x)}{u}\right)-g^{2}(x)+u\right)\,.$
In particular, denoting by $\mu$ the uniform measure over $\Omega^{*}$ and taking $u=\mathbb{E}_{\mu}[g^{2}]$, we have
$\operatorname{Ent}_{\pi}(f^{2})$ $\leq$ $\sum_{x\in\Omega^{*}}\frac{1}{|\Omega|}\left(g^{2}(x)\log\left(\frac{g^{2}(x)}{\mathbb{E}_{\mu}[g^{2}]}\right)-g^{2}(x)+\mathbb{E}_{\mu}[g^{2}]\right)$
$=$ $\frac{2^{n^{2}}}{|\Omega|}\operatorname{Ent}_{\mu}(g^{2})\,.$
If $X$ is a random matrix with distribution $\mu$, the rows $L_{1},\dots,L_{n}$ of $X$ are independent, so by the sub-additivity of entropy for product measures, we have
$\operatorname{Ent}_{\mu}(g^{2})$ $\leq$ $\sum_{i=1}^{n}\mathbb{E}_{\mu}\left[\operatorname{Ent}_{\mu}\left(g^{2}\,|\,(L_{k})_{k\neq i}\right)\right]$
For each $i\in[n]$, conditionally on $(L_{k})_{k\neq i}=(\ell_{k})_{k\neq i}$, observe that if the family $(\ell_{k})_{k\neq i}$ does not span a subset of dimension n1n-1, then, whatever the value of $L_{i}$, the function $g$ is equal to $\mathbb{E}_{\pi}[f]$, so the conditional entropy is zero. Otherwise, if the family $(\ell_{k})_{k\neq i}$ spans a subset of dimension n1n-1, then one may complete it with some vector $\ell_{i}$ to form a basis of $\{0,1\}^{n}$. The conditional law of $L_{i}$ is then that of $\sum_{k=1}^{n}a_{k}\ell_{k}$, where $(a_{k})_{k=1}^{n}$ is uniformly distributed over $\{0,1\}^{n}$. Since the simple random walk on the hypercube has logarithmic Sobolev constant equal to $n$ (see for instance [4, Example 3.2]), we have, for each $i\in[n]$,
$\operatorname{Ent}_{\mu}\left(g^{2}\,|\,(L_{k})_{k\neq i}=(\ell_{k})_{k\neq i}\right)\leq\frac{1}{2}\sum_{j=1}^{n}\mathbb{E}_{\mu}\left[\left(g(X)-g(X^{i\leftarrow j})\right)^{2}\,\big|\,(L_{k})_{k\neq i}=(\ell_{k})_{k\neq i}\right]\,$,
@@
One may now conclude with the results of [7, Section 3.2] on the spectral gap, which yield
$\operatorname{Var}_{\pi}(f)\;\leq\;4(31\sqrt{n}+700)^{2}{\mathcal{E}}_{P}(f,f)\,.$